The sum of two numbers is 30. Find the maximum possible product of the two numbers.
Any help would be greatly appreaciated
Thank you ^_^

Answer by JulietG(1812)

The maximum product is going to be the two numbers that are closest together.
15 + 15 = 30
15 * 15 = 225
If the numbers need to be different, then 14 & 16
14 + 16 = 30
14 * 16 = 224